OpenStudy (anonymous):

review of alg 2 helpp?

OpenStudy (abb0t):

Seeking a review or asking for one?

OpenStudy (anonymous):

\[\frac{ 125 }{ 8 } \]

OpenStudy (anonymous):

to the -2/3

OpenStudy (abb0t):

if you take the cube root, you can get \(\frac{5}{2}\)

OpenStudy (anonymous):

the answer is 4/25 but thats what i thought too

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

cuz it''s 2/3

OpenStudy (agent0smith):

\[\Large \left( \frac{ 125 }{ 8 } \right)^{-2/3} = \left( \frac{ 8 }{ 125 } \right)^{2/3} \] \[\Large \left( \frac{ 8^{1/3} }{ 125^{1/3} } \right)^{2} = \left( \frac{ 2 }{ 5 } \right)^2\] I'm sure you can finish.